OpenStudy (anonymous):

A farmer was counting his cows and chickens he counted 13 heads and 36 legs how many cows does he have

OpenStudy (shamim):

U know both cow nd hen has only 1 head

OpenStudy (shamim):

But every cow has 4 legs

OpenStudy (shamim):

Every hen has 2 legs

OpenStudy (shamim):

Farmer counts 13 heads

OpenStudy (shamim):

So total number of cows nd hens =13

OpenStudy (shamim):

Number of cows+number of hens=13 Right?

OpenStudy (shamim):

Now let the number of cows=x

OpenStudy (shamim):

So the number of hens=13-x

OpenStudy (shamim):

So total legs of all cows=4x

OpenStudy (shamim):

Total legs of all hens=(13-x)*2

OpenStudy (shamim):

So total legs of cows nd hens=4x+(13-x)*2=36

OpenStudy (shamim):

x=?

OpenStudy (shamim):

Response plz for better understanding

OpenStudy (shubhipapa):

let no of cows=x cow has 4 legs so legs of cow=4x no of chikens=13-x no of legs of chicken=2(13-x) total no of legs=36 so 4x+2(13-x)=36 4x+26-2x=36 2x=10 x=5 so no of cows=5 no of chickens=13-5= 8